# RISE: 3D Perception Makes Real-World Robot Imitation Simple and Effective

### v2

@@ I Introduction
Spatial information can also be represented using 3D point clouds, which have demonstrated remarkable advantages in robotic tasks like general grasping [anygrasp, sundermeyer2021contact]. Recently, imitation learning based on point clouds is drawing increasing interest in our community [chen2023polarnet, gervet23act3d, goyal2023rvt, guhur2022hiveformer, james2022coarse, shridar2022peract, xian2023chaineddiffuser, ze2023gnfactor]. Most of the 3D-based methods learn to predict the next keyframe as opposed to continuous actions, which often struggle with tasks involving frequent contacts and abrupt environmental changes. Meanwhile, addressing the annotation of keyframes at scale for real-world data necessitates additional manual effort. The reason that they struggle to predict continuous actions lies in the difficulty of achieving real-time performance, which relies heavily on efficient 3D feature encoding and remains a challenge for point cloud based methods.
@@ II-A Imitation Learning for Robotics
3D Imitation Learning. The formulation of incorporating 3D information in the imitation learning framework is under active exploration. The most straightforward method is to apply projections to transform the 3D point cloud to several 2D image views and transfer the task to multi-view image-based policy learning [goyal2023rvt, guhur2022hiveformer], which requires the virtual viewpoints to be carefully designed to ensure performance. Moreover, due to sparse and noisily sensed point clouds, [goyal2023rvt] fails to grasp slim objects like marker pens in real-world experiments. [james2022coarse, shridar2022peract, ze2023gnfactor] process point clouds to dense voxel grids and apply 3D convolutions. Since high-resolution 3D feature maps require expensive computes, these methods have to trade off performance against cost. [gervet23act3d, xian2023chaineddiffuser] featurize point clouds by projecting multi-view image features to 3D world to avoid dense convolutions. However, such feature fusion techniques struggle to capture the consistent 3D representation from different views accurately. Recently, a concurrent work DP3 [ze2024dp3] also leverages 3D perception in robotic manipulation policies, but our real-world evaluations in §IV-F demonstrate that it cannot handle demonstrations with various representations limited by its network capacity.
@@ IV-B Pick-and-Place Tasks
TABLE I: Experimental results of the 6-DoF task Pour Balls.
@@ IV-F Effectiveness of 3D Perception
We also evaluate the recently proposed DP3 [ze2024dp3] in this experimental setting. However, DP3 appears to struggle in learning meaningful actions from our demonstration data. After communicating with the authors, one potential reason is that they are using RealSense L515 in their original experiments, and we adopt RealSense D435 in our experiments. The point cloud from D435 is noiser, making it more challenging for networks to learn. By using sparse convolution, RISE is more robust to the noise in point cloud. Besides, after delving into their real robot experiments, we found that instead of natural actions, axis-wise actions are used in their demonstration data, as illustrated in Tab. V (left). Hence, we collect 50 demonstrations on the Collect Cups task with 1 cup using axis-wise and natural action representations respectively. These demonstrations are then used for DP3 policy training. After carefully tuning hyper-parameters such as horizons and color utilizations, we report the evaluation results in Tab. V, with the best completion rate of 40%. We suspect that the limited network capacity of the 3D encoder of DP3 prevents it from modeling the diverse state-action pairs present in the real-world human teleoperated demonstrations, leading it to only handle a smaller set of state-action pairs under the axis-wise action representations. On the contrary, RISE can handle real-world demonstrations with various action representations and maintain satisfactory performances. Lastly, compared to the evaluation setup in the DP3 paper, we allow objects to be placed anywhere in the entire workspace. This results in a greater variation of object locations, making the task more challenging.
@@ V-B Implementation Details
V-C Discussions about Action Representations
@@
V-D Failure Cases and Recovery

### v3

@@ I Introduction
Spatial information can also be represented using 3D point clouds, which have demonstrated remarkable advantages in robotic tasks like general grasping [anygrasp, sundermeyer2021contact]. Recently, imitation learning based on point clouds is drawing increasing interest in our community [chen2023polarnet, gervet23act3d, goyal2023rvt, guhur2022hiveformer, james2022coarse, shridar2022peract, xian2023chaineddiffuser, ze2023gnfactor, sgr23]. Most of the 3D-based methods learn to predict the next keyframe as opposed to continuous actions, which often struggle with tasks involving frequent contacts and abrupt environmental changes. Meanwhile, addressing the annotation of keyframes at scale for real-world data necessitates additional manual effort. The reason that they struggle to predict continuous actions lies in the difficulty of achieving real-time performance, which relies heavily on efficient 3D feature encoding and remains a challenge for point cloud based methods.
@@ II-A Imitation Learning for Robotics
3D Imitation Learning. The formulation of incorporating 3D information in the imitation learning framework is under active exploration. The most straightforward method is to apply projections to transform the 3D point cloud to several 2D image views and transfer the task to multi-view image-based policy learning [goyal2023rvt, guhur2022hiveformer], which requires the virtual viewpoints to be carefully designed to ensure performance. Moreover, due to sparse and noisily sensed point clouds, [goyal2023rvt] fails to grasp slim objects like marker pens in real-world experiments. [james2022coarse, shridar2022peract, ze2023gnfactor] process point clouds to dense voxel grids and apply 3D convolutions. Since high-resolution 3D feature maps require expensive computes, these methods have to trade off performance against cost. [zhu2023learning] proposed an object-centric representation for learning which requires an additional segmentation process. [gervet23act3d, xian2023chaineddiffuser] featurize point clouds by projecting multi-view image features to 3D world to avoid dense convolutions. However, such feature fusion techniques struggle to capture the consistent 3D representation from different views accurately. Recently, a concurrent work DP3 [ze2024dp3] also leverages 3D perception in robotic manipulation policies, but our real-world evaluations in §IV-F demonstrate that it cannot handle demonstrations with various representations limited by its network capacity. DexCap [wang2024dexcap] also proposed a PointNet [qi2017pointnet] with diffusion head architecture for dexterous manipulation.
@@ IV-B Pick-and-Place Tasks
TABLE I: Experimental results of the 6-DoF task Pour Balls. “If Poured” denotes the completion rate in the context of having completed the pouring action.
@@ IV-F Effectiveness of 3D Perception
We also evaluate the recently proposed DP3 [ze2024dp3] in this experimental setting. However, DP3 appears to struggle in learning meaningful actions from our demonstration data. After communicating with the authors, one potential reason is that they are using RealSense L515 in their original experiments, and we adopted RealSense D435 in our experiments. The point cloud from D435 is noisier, making it more challenging for networks to learn. By using sparse convolution, RISE is more robust to the noise in the point cloud. Besides, after delving into their real robot experiments, we found that instead of natural actions, axis-wise actions are used in their demonstration data, as illustrated in Tab. V (left). Hence, we collect 50 demonstrations on the Collect Cups task with 1 cup using axis-wise and natural action representations respectively. These demonstrations are then used for DP3 policy training. After carefully tuning hyper-parameters such as horizons and color utilizations, we report the evaluation results in Tab. V, with the best completion rate of 40%. We suspect that the limited network capacity of the 3D encoder of DP3 prevents it from modeling the diverse state-action pairs present in the real-world human teleoperated demonstrations, leading it to only handle a smaller set of state-action pairs under the axis-wise action representations. On the contrary, RISE can handle real-world demonstrations with various action representations and maintain satisfactory performances. Lastly, compared to the evaluation setup in the DP3 paper, we allow objects to be placed anywhere in the entire workspace. This results in a greater variation of object locations, making the task more challenging.
@@ V Conclusion
VI Acknowledgement
This work was supported by the National Key Research and Development Project of China (No. 2022ZD0160102), the National Key Research and Development Project of China (No. 2021ZD0110704), Shanghai Artificial Intelligence Laboratory, XPLORER PRIZE grants.
@@ VI-B Implementation Details
VI-C Discussions about Action Representations
@@
VI-D Failure Cases and Recovery